# T-LoRA: Single Image Diffusion Model Customization Without Overfitting

### v1

@@ 3.4 T-LoRA
Finally, we present the complete T-LoRA framework (Figure 2(c)), which integrates the timestep-dependent rank control of Vanilla T-LoRA with the orthogonality-enforcing initialization of Ortho-LoRA, yielding a timestep adaptive solution for diffusion model personalization.:
@@
where $A_{init}=V^{T}[-r:]$, $B_{init}=U[-r:]$ and $S_{init}=S[-r:]$ are initiaized with the last SVD components of a random matrics $R=USV^{T},R\sim\mathcal{N}(0,1/r)$. And $r(t)=\lfloor(r-r_{min})\cdot(T-t)/T\rfloor+r_{min}$ as timesteps progress towards lower values.
@@ 4 Experiments
Dataset We use $25$ concepts from prior works (TI, ; DB, ; CD, ), covering a diverse range of categories including pets, toys, interior objects, accessories, and more. Only one image per concept is used for training. This images are manually selected to ensure clear concept visibility. Each concept is paired with $25$ text prompts providing appearance, position, and background modifications, and 6 complex prompts involving multiple modifications (e.g., background and accessorization). A full list of prompts is provided in Appendix A. For each concept, we generate $5$ images per contextual prompt and $15$ per base prompt (e.g., "a photo of $V^{*}$"), totaling in 800 unique concept-prompt pairs. All experiments, except for the Initialization Investigation, involve full 25 concepts, for Initialization Investigation we randomly sampled 5 concepts to reduce computational demands.
Evaluation Metric To assess concept fidelity, we calculate the average pairwise cosine similarity (IS) between CLIP ViTB/32 clip embeddings of real and generated images, as described in TI . This approach allows us to maintain the measurement’s independence from positional, appearance, and accessorization modifications by using the scene-independent prompt "a photo of $V^{*}$." Higher values of this metric typically indicate better subject fidelity. To further eliminate the influence of the training images background reproduction, we mask the background. Additionally, we provide DINO-IS, which is calculated the same way, but using DINO dino embeddings. To evaluate the correspondence between generated images and contextual prompts (TS), we compute the average cosine similarity between the CLIP ViTB/32 embeddings of the prompt and the generated images.
Experimental Setup In all our experiments, we train the Stable Diffusion-XL sdxl model with a batch size of 1, fine-tuning only the diffusion U-Net while keeping the text encoder fixed. Baseline models follow their original paper setups. For additional training details, please refer to Appendix B.
@@ 4.1 T-LoRA Design Decisions
Analysis of Ortho-LoRA Initialization Strategies We investigated six variants of the Ortho-LoRA initialization, which are based on the principal, middle, and last components of the original weights $W$, and a random matrix $R$. The IS and TS metrics for these setups are presented in Figure 5. First, we observe that for most ranks, the points in the TS metric are ordered according to the initialization singular values: the principal components initialization yields the lowest TS, followed by the middle components initialization, and the last components initialization yields the highest TS. This suggests that higher singular values are strongly correlated with overfitting, particularly for the $W$ SVD initialization, which has the largest principal singular values (see Figure 6). Second, for ranks 4 and 8, we observe that initializing with the last SVD components of $W$ slows down the training process, whereas initializing with the last SVD components of $R$ does not have this effect. This discrepancy is likely due to the smaller values produced by the last SVD components of $W$ compared to those initialized with a random matrix $R$ (as illustrated in the close-up image of the last singular values in Figure 6). Overall, initializing with the last SVD components from a random matrix $R$ yields optimal results for most ranks, that is why we use it in all further experiments with Ortho-LoRA and T-LoRA.
Figure 6: Comparison between singular values distribution from the model’s self-attention layers (size 1280) and from a random matrix $R$ of the same size, with a close-up on the smallest singular values.
Figure 7: Generation outputs for Ortho-LoRA and T-LoRA with $r_{min}$ set to 25% and 50% of the full rank $r=64$. Ortho-LoRA generally exhibits poor alignment with the text and a high degree of reproduction of training images. In contrast, T-LoRA significantly enhances alignment with the text. Notably, T-LoRA at 50% maintains high image fidelity, whereas T-LoRA at 25% often struggles to accurately reproduce the concept.
4.2 Comparison with LoRA
@@
In Table 1, we present image and text similarity for LoRA, Vanilla T-LoRA, and T-LoRA across various ranks. For all ranks, both Vanilla T-LoRA and T-LoRA demonstrate superior text similarity compared to LoRA, while maintaining high image similarity that differs from LoRA by only a third of a decimal place. At lower ranks, Vanilla T-LoRA and T-LoRA show similar performance; however, the performance improvement of T-LoRA becomes more pronounced as the rank increases. At low ranks, LoRA approaches full rank, which is why Vanilla T-LoRA performs comparably to T-LoRA. In contrast, as the ranks increase, the effectiveness of masking in Vanilla T-LoRA diminishes, while T-LoRA continues to demonstrate its full potential.
4.3 Comparison with Baselines
@@
Metric T-LoRA-64 LoRA-64 lora OFT-32 ortogonal OFT-16 ortogonal GSOFT-64 groupshouffle GSOFT-32 groupshouffle SVDiff svdiff
@@
In addition to LoRA lora , we compare our T-LoRA with other lightweight customization methods, including OFT ortogonal , GSOFT groupshouffle , and SVDiff svdiff . The results are presented in Table 2.
Our T-LoRA achieves the best text similarity across all methods, except for SVDiff; however, SVDiff exhibits very low image similarity and often fails to accurately represent the concept. While LoRA demonstrates the highest image similarity, it also exhibits the most significant overfitting. Notably, our method’s image similarity differs from LoRA’s by only a third of a decimal place.
Figure 8 showcases examples of generation for each method. T-LoRA provides greater flexibility in generation concerning position and background changes while accurately representing the concept.
Figure 8: Generation examples for T-LoRA alongside other diffusion model customization baselines.
4.4 Multi-image Experiments
In addition to the single-image experiments presented in the main paper, we evaluate T-LoRA against LoRA lora and OFT ortogonal in the context of multi-image diffusion model customization. In this setting, each concept is represented by multiple images featuring diverse backgrounds. Table 3 summarizes the results for experiments conducted with 1, 2, and 3 images. For T-LoRA and LoRA we use $r=64$, and $r=32$ for OFT as it showed the best results in single-image setup.
Our results demonstrate that T-LoRA consistently outperforms LoRA in terms of text similarity across all numbers of images, while achieving comparable image similarity. Notably, T-LoRA trained on a single image surpasses LoRA trained on 2 or even 3 images of the same concept. When compared to OFT, T-LoRA achieves superior results in the 2-image scenario, while their performances are comparable in the 3-image scenario.
Table 3: Image Similarity (IS) and Text Similarity (TS) for multi-image customization setting.
Methods # Images = 1 # Images = 2 # Images = 3
IS TS IS TS IS TS
LoRA-64 lora 0.901 0.232 0.900 0.245 0.902 0.251
OFT-32 ortogonal 0.901 0.247 0.901 0.261 0.901 0.267
T-LoRA-64 0.900 0.256 0.901 0.262 0.900 0.263
4.5 User Study
@@
LoRA-64 lora 39.3 60.7 71.0 29.0 67.3 32.7
OFT-32 ortogonal 52.5 47.5 58.3 41.7 63.5 36.5
GSOFT-64 groupshouffle 49.0 51.0 61.5 38.5 60.3 39.7
SVDiff svdiff 90.1 9.9 42.1 57.9 55.9 44.1
@@
Finally, we conduct Human Evaluation to fully investigate our model’s performance. Using an original image of the concept, a text prompt, and two generated images (one from T-LoRA and the other from an alternative method), we asked users to respond to the following questions: (1) "Which image more accurately represents the original concept?" to evaluate image similarity (2) "Which image aligns more closely with the text prompt?" to assess text similarity; and (3) "Which image overall demonstrates better alignment with the prompt and preserves the identity of the concept?" to evaluate the overall preference. For each pair of methods, we randomly generated 60 unique concept-prompt pairs. In total, we collect 1,800 human assessments across six method pairs.
@@
Results in Table 4 show that T-LoRA significantly outperforms others in text similarity and overall preference, while achieving comparable or superior image similarity to most methods. The exception is LoRA, which surpasses T-LoRA in image similarity due to its tendency to overfit and fully reproduce the original object. Despite this, T-LoRA maintains a strong overall impression, highlighting its balanced performance across criteria.
@@ 5 Conclusion
This paper addressed the challenge of personalizing diffusion models using a single concept image, where overfitting and limited generative diversity are prevalent. We introduced T-LoRA, a Timestep-Dependent Low-Rank Adaptation framework, featuring (1) a rank masking strategy to regulate training across diffusion timesteps and (2) Ortho-LoRA, an orthogonal weight initialization technique to enhance effective rank utilization. Extensive experiments demonstrated that T-LoRA significantly outperforms existing approaches, achieving a balance between concept fidelity and text alignment in data-limited settings. These findings lay a strong foundation for integrating timestep-sensitive adaptation strategies and orthogonality principles into future diffusion model frameworks, with promising implications for creative applications in text-to-image generation and beyond.
Appendix A Prompts
@@
live_set = [ a chocolate-colored V*’, a red V*’, a purple V*’, a brightly lit V*’, a snowy V*’, a V* in a purple wizard outfit’, a V* in a doctor outfit’, a V* in a superhero suit’, a V* dressed as a ballerina’, a V* wearing a green bow tie’, a V* with the Eiffel Tower in the background’, a V* with a mountain in the background’, a V* with a tree and autumn leaves in the background’, a V* on top of green grass with sunflowers around it’, a V* with a sunset and palm trees in the background’, a V* with the Sydney Opera House in the background’, a V* lying in the bed’, a V* sitting on the chair’, a V* swimming ’, a V* sleeping ’, a V* skateboarding’, a V* riding a surfboard’, a V* jumping with a parachute’, a V* running on a treadmill’, a V* riding a bike’, a V* in a chief outfit in a nostalgic kitchen filled with vintage furniture and scattered biscuit’, a V* dressed in a superhero cape, soaring through the skies above a bustling city during a sunset’, a V* wearing a tuxedo performing on a stage in the spotlight’, "a V* in a wizards robe in a magical forest at midnight, accented with purples and sparkling silver tones", a V* wearing a space helmet, among stars in a cosmic landscape during a starry night’, a V* wearing a pirate hat exploring a sandy beach at the sunset with a boat floating in the background’, ] object_set = [ a chocolate-colored V*’, a V* made of gold’, a brightly lit V*’, a red V*’, a purple V*’, a V* on top of a red and white checkered picnic blanket’, a V* with the Eiffel Tower in the background’, a V* with a mountain in the background’, a V* with a tree and autumn leaves in the background’, a V* on top of green grass with sunflowers around it’, a V* with a sunset and palm trees in the background’, a V* with the Sydney Opera House in the background’, a V* in the snow’, a V* on the beach’, a V* on a cobblestone street’, a V* on top of a wooden floor’, a V* with a city in the background’, a V* with a blue house in the background’, a V* on top of a purple rug in a forest’, a V* with a wheat field in the background’, a V* in the bed’, a V* on the chair’, a V* on top of a snowy mountain peak’, a V* with a sand castle in the background’, a V* in a cave’, a V* on a windowsill in Tokyo at dusk, illuminated by neon city lights, using neon color palette’, a V* on a sofa in a cozy living room, rendered in warm tones’, a V* on a wooden table in a sunny backyard, surrounded by flowers and butterflies’, a V* floating in a bathtub filled with bubbles and illuminated by the warm glow of evening sunlight filtering through a nearby window’, a V* floating among stars in a cosmic landscape during a starry night with a spacecraft in the background’, a V* on a sandy beach next to the sand castle at the sunset with a floaing boat in the background’, ]
Appendix B Training Details
[…]
OFT We implement OFT with PEFT library [25]. We train all OFT models for 800 training steps.
GSOFT We use offitial repo https://github.com/Skonor/group_and_shuffle for GSOFT implementation. We train all GSOFT models for 800 training steps.
[…]
Appendix C Limitations

### v2

@@ 3.3 T-LoRA
Finally, we introduce the complete T-LoRA framework (Figure 3(c)), combining Vanilla T-LoRA’s timestep-dependent rank control with Ortho-LoRA’s orthogonal initialization, resulting in a timestep-adaptive solution for diffusion model personalization:
@@
where $A_{init}=V^{T}[-r:]$, $B_{init}=U[-r:]$ and $S_{init}=S[-r:]$ are the last SVD components of a random matrix $R=USV^{T},R\sim\mathcal{N}(0,1/r)$. The rank schedule follows $r(t)=\lfloor(r-r_{min})\cdot(T-t)/T\rfloor+r_{min}$. We do not apply any orthogonality-enforcing regularization, as the Ortho-LoRA initialization inherently maintains the orthogonality of matrices throughout the entire training process.
@@ 4 Experiments
Dataset We use 25 concepts from prior works (Gal et al. 2022; Ruiz et al. 2023; Kumari et al. 2023), including pets, toys, interior objects, accessories, and more. Each concept is trained on a single, manually selected image with clear visibility. For evaluation, each is paired with 25 contextual prompts (appearance, position, background changes) and 6 complex prompts (e.g., background + accessorization). We generate 5 images per contextual prompt and 15 per base prompt (e.g., ”a photo of $V^{*}$”), totaling 800 concept-prompt pairs. All experiments use all 25 concepts, except the Initialization Investigation, which uses 5 randomly sampled concepts to reduce compute.
Evaluation Metric To assess concept fidelity, we compute the average pairwise cosine similarity (IS) between CLIP ViT-B/32 (Radford et al. 2021) embeddings of real and generated images, following (Gal et al. 2022). Using the neutral prompt ”a photo of $V^{*}$” ensures independence from appearance, position, and accessorization changes. Backgrounds are masked to reduce bias from training image reproduction. We also report DINO-IS, computed similarly with DINO (Caron et al. 2021) embeddings. To evaluate prompt-image alignment (TS), we calculate the average cosine similarity between CLIP embeddings of the prompt and generated images.
Experimental Setup We investigated the performance of our method on two models: Stable Diffusion-XL (Podell et al. 2023) and FLUX-1.dev (Labs 2024). In all experiments models are fine-tuned with batch size 1, updating only the diffusion U-Net/DiT while keeping the text encoder fixed. Baselines follow their original setups.
Figure 6: Ablation study. Generation outputs for Ortho-LoRA and T-LoRA with $r_{min}$ set to 25% and 50% of the full rank $r=64$. Ortho-LoRA generally exhibits poor alignment with the text and a high degree of reproduction of training images. In contrast, T-LoRA significantly enhances alignment with the text. Notably, T-LoRA at 50% maintains high image fidelity, whereas T-LoRA at 25% often struggles to accurately reproduce the concept.
@@ 4.1 T-LoRA Design Decisions
Metric T-LoRA-64 LoRA-64 OFT-32 OFT-16 GSOFT-64 GSOFT-32 SVDiff
@@
LoRA-64 39.3 60.7 71.0 29.0 67.3 32.7
OFT-32 52.5 47.5 58.3 41.7 63.5 36.5
GSOFT-64 49.0 51.0 61.5 38.5 60.3 39.7
SVDiff 90.1 9.9 42.1 57.9 55.9 44.1
Analysis of Ortho-LoRA Initialization Strategies We investigated six variants of the Ortho-LoRA initialization, which are based on the principal, middle, and last components of the original weights $W$, and a random matrix $R$. The IS and TS metrics for these setups are presented in Figure 5. First, we observe that for all ranks and for both $R$ and $W$ initializations the points in the TS metric are ordered according to the initialization singular values magnitude: the principal components initialization yields the lowest TS, followed by the middle components initialization, and the last components initialization yields the highest TS. This suggests that higher singular values are strongly correlated with overfitting. For ranks 4 and 8, initializing with the last SVD components of $W$ turns out to be too close to zero and slows down the training process, whereas initializing with the last SVD components of $R$ does not have this effect. Overall, initializing with the last SVD components from a random matrix $R$ yields optimal results for most ranks, that is why we use it in all further experiments with Ortho-LoRA and T-LoRA.
4.2 Comparison with LoRA
In Table 1, we present image and text similarity for LoRA, Vanilla T-LoRA, and T-LoRA across various ranks on SD-XL. For all ranks, both Vanilla T-LoRA and T-LoRA demonstrate superior text similarity compared to LoRA, while maintaining same image similarity that differs from LoRA by only a third of a decimal place. At lower ranks, Vanilla T-LoRA and T-LoRA show similar performance; however, the performance improvement of T-LoRA becomes more pronounced as the rank increases. At low ranks, LoRA approaches full rank, which is why Vanilla T-LoRA performs comparably to T-LoRA. In contrast, as the ranks increase, the effectiveness of masking in Vanilla T-LoRA diminishes, while T-LoRA continues to demonstrate its full potential.
4.3 Comparison with Baselines
In addition to LoRA (Hu et al. 2021), we compare our T-LoRA with other lightweight customization methods, including OFT (Qiu et al. 2023), GSOFT (Gorbunov et al. 2024), and SVDiff (Han et al. 2023). The results on SD-XL are presented in Table 2.
T-LoRA achieves the best text similarity across all methods, except for SVDiff; however, SVDiff exhibits very low image similarity and often fails to accurately represent the concept. While LoRA demonstrates the highest image similarity, it also exhibits the most significant overfitting. Notably, our method’s image similarity differs from LoRA’s by only a third of a decimal place.
Figure LABEL:fig:all_methods showcases examples of generation for each method. T-LoRA provides greater flexibility in generation concerning position and background changes while accurately representing the concept.
4.4 Multi-image Experiments
In addition to the single-image experiments, we evaluate T-LoRA against LoRA (Hu et al. 2021) and OFT (Qiu et al. 2023) in the multi-image diffusion model customization. In this setting, each concept is represented by multiple images featuring diverse backgrounds. Table 4 summarizes the results for experiments on SD-XL conducted with 1, 2, and 3 images. For T-LoRA and LoRA we use $r=64$, and $n_{blocks}=32$ for OFT as it showed the best results in single-image setup.
T-LoRA consistently outperforms LoRA in text similarity across all image counts while achieving similar image similarity. Remarkably, T-LoRA trained on one image surpasses LoRA trained on two or three images. Compared to OFT, T-LoRA excels in the two-image scenario and performs similarly in the three-image scenario.
Methods # Images = 1 # Images = 2 # Images = 3
IS TS IS TS IS TS
LoRA-64 0.901 0.232 0.900 0.245 0.902 0.251
OFT-32 0.901 0.247 0.901 0.261 0.901 0.267
T-LoRA-64 0.900 0.256 0.901 0.262 0.900 0.263
Table 4: Image Similarity (IS) and Text Similarity (TS) for multi-image Customization experiments.
4.5 User Study
@@
We also conduct Human Evaluation to fully investigate our model’s performance. Using an original image of the concept, a text prompt, and two generated images (one from T-LoRA and the other from an alternative method), we asked users to respond to the following questions: (1) ”Which image more accurately represents the original concept?” to evaluate image similarity (2) ”Which image aligns more closely with the text prompt?” to assess text similarity; and (3) ”Which image overall demonstrates better alignment with the prompt and preserves the identity of the concept?” to evaluate the overall preference. For each pair of methods, we randomly generated 60 unique concept-prompt pairs. In total, we collect 1,800 human assessments across six method pairs.
@@
Results in Table 3 show that T-LoRA significantly outperforms others in text similarity and overall preference, while achieving comparable or superior image similarity to most methods. The exception is LoRA, which surpasses T-LoRA in image similarity due to its tendency to overfit and fully reproduce the original image. Despite this, T-LoRA maintains a strong overall impression, highlighting its balanced performance across criteria.
4.6 FLUX-1.dev Experiments
Finally, we investigated the performance of our model on the MM-DiT flow-based model, using FLUX-1.dev (Labs 2024). For these experiments, we randomly selected 10 concepts from the full set: five live concepts and five objects. The prompt set was the same as in the SD-XL experiments.
We found that, unlike SD-XL, LoRA adapters trained on MM-DiT are full-rank (see Figure 9), making the additional orthogonal initialization unnecessary. Therefore, we used the Vanilla variant of T-LoRA for all FLUX-1.dev experiments.
The results are presented in Table 5 and Figure LABEL:fig:flux_all. T-LoRA consistently achieves better prompt alignment and produces images with a more consistent ambiance than LoRA, as confirmed by both quantitative metrics and qualitative visual examples.
Figure 9: Singular values of FLUX-1.dev LoRA B matrices with $r=64$ after 500 training steps. The red line marks the rank capturing 95% of the total singular value sum. Compared to SD-XL, all the singular values are non-zero and effective rank is near the full rank.
Table 5: Image Similarity (IS) and Text Similarity (TS) for LoRA and T-LoRA across dfferent ranks on FLUX-1.-dev.
Methods Rank = 4 Rank = 16 Rank = 64
IS TS IS TS IS TS
LoRA 0.890 0.263 0.905 0.264 0.884 0.247
T-LoRA 0.908 0.268 0.903 0.280 0.888 0.280
@@ 5 Conclusion
This paper addressed the challenge of personalizing diffusion models using a single concept image, where overfitting and limited generative diversity are prevalent. We introduced T-LoRA, a Timestep-Dependent Low-Rank Adaptation framework featuring (1) a rank masking strategy to regulate training across diffusion timesteps and (2) Ortho-LoRA, an orthogonal weight initialization technique to enhance effective rank utilization. Extensive experiments show that T-LoRA outperforms prior methods, balancing concept fidelity and text alignment in data-limited settings. Our results highlight the potential of timestep-aware adaptation and orthogonality for advancing text-to-image generation and related creative tasks.
Appendix A Limitations
Appendix B Training Details
B.1 SD-XL
All the models, except SVDiff are trained using Adam optimizer with $\text{batch size}=1$, $\text{learning rate}=1e-4$, $\text{betas}=(0.9,0.999)$ and $\text{weight decay}=1e-4$. All the experiments were conducted using a single GPU H-100 per each model customization.
OFT We implement OFT with PEFT library (Mangrulkar et al. 2022). We train all OFT models for 800 training steps.
GSOFT We use offitial repo https://github.com/Skonor/group˙and˙shuffle for GSOFT implementation. We train all GSOFT models for 800 training steps.
SVDiff We implement the method based on https://github.com/mkshing/svdiff-pytorch. The model for all concepts were trained for $1600$ training steps using Adam optimizer with $\text{batch size}=1$, $\text{learning rate}=1e-3$, $\text{learning rate 1d}=1e-6$, $\text{betas}=(0.9,0.999)$, $\text{epsilon}=1e\!-\!8$, and $\text{weight decay}=0.01$.
B.2 FLUX-1.dev
All the models are trained using prodigy optimizer with $\text{batch size}=1$, $\text{learning rate}=1$. We trained LoRA and T-LoRA for $400$ and $500$ training steps, respectively.
Figure 10: Generation examples for LoRA and T-LoRA $r=64$ on the different training steps.
Figure 11: Generation examples for LoRA and T-LoRA $r=16$ on the different training steps.
Figure 12: Generation examples for LoRA and T-LoRA $r=4$ on the different training steps.
Appendix C Training Dynamics
In Figures 10, 11, and 12, we present generation examples for LoRA and T-LoRA at different training steps across various ranks ($r=[64,16,4]$). A common observation during LoRA fine-tuning is that it tends to overfit, producing examples with poor prompt alignment while it has not yet fully learned the concept. This effect is particularly evident in the 200 training steps for the duck toy (Figure 10), where the third image lacks an appropriate background, while the first two images fail to accurately represent the concept. Similarly, in the example with the cat (Figure 10), the generation at 200 steps includes a flower from the training set, while the cat itself is not represented correctly. In contrast, T-LoRA typically requires slightly longer training to grasp the concept, but it learns more smoothly and exhibits significantly less overfitting in terms of both position and background. This observation is consistent across the other ranks as well.
Appendix D Prompts
@@
live_set = [ a chocolate-colored V*’, a red V*’, a purple V*’, a brightly lit V*’, a snowy V*’, a V* in a purple wizard outfit’, a V* in a doctor outfit’, a V* in a superhero suit’, a V* dressed as a ballerina’, a V* wearing a green bow tie’, a V* with the Eiffel Tower in the background’, a V* with a mountain in the background’, a V* with a tree and autumn leaves in the background’, a V* on top of green grass with sunflowers around it’, a V* with a sunset and palm trees in the background’, a V* with the Sydney Opera House in the background’, a V* lying in the bed’, a V* sitting on the chair’, a V* swimming ’, a V* sleeping ’, a V* skateboarding’, a V* riding a surfboard’, a V* jumping with a parachute’, a V* running on a treadmill’, a V* riding a bike’, a V* in a chief outfit in a nostalgic kitchen filled with vintage furniture and scattered biscuit’, a V* dressed in a superhero cape, soaring through the skies above a bustling city during a sunset’, a V* dressed wearing a tuxedo performing on a stage in the spotlight’, "a V* in a wizards robe in a magical forest at midnight, accented with purples and sparkling silver tones", a V* wearing a space helmet, among stars in a cosmic landscape during a starry night’, a V* wearing a pirate hat exploring a sandy beach at the sunset with a boat floating in the background’, ] object_set = [ a chocolate-colored V*’, a V* made of gold’, a brightly lit V*’, a red V*’, a purple V*’, a V* on top of a red and white checkered picnic blanket’, a V* with the Eiffel Tower in the background’, a V* with a mountain in the background’, a V* with a tree and autumn leaves in the background’, a V* on top of green grass with sunflowers around it’, a V* with a sunset and palm trees in the background’, a V* with the Sydney Opera House in the background’, a V* in the snow’, a V* on the beach’, a V* on a cobblestone street’, a V* on top of a wooden floor’, a V* with a city in the background’, a V* with a blue house in the background’, a V* on top of a purple rug in a forest’, a V* with a wheat field in the background’, a V* in the bed’, a V* on the chair’, a V* on top of a snowy mountain peak’, a V* with a sand castle in the background’, a V* in a cave’, a V* on a windowsill in Tokyo at dusk, illuminated by neon city lights, using neon color palette’, a V* on a sofa in a cozy living room, rendered in warm tones’, a V* on a wooden table in a sunny backyard, surrounded by flowers and butterflies’, a V* floating in a bathtub filled with bubbles and illuminated by the warm glow of evening sunlight filtering through a nearby window’, a V* floating among stars in a cosmic landscape during a starry night with a spacecraft in the background’, a V* on a sandy beach next to the sand castle at the sunset with a floaing boat in the background’, ]